OpenStudy (anonymous):

solve for y: x=(y+8)/(y+4)

OpenStudy (anonymous):

\[x=\left(\begin{matrix}y+8 \\ y+4\end{matrix}\right)\]

jimthompson5910 (jim_thompson5910):

x=(y+8)/(y+4) x(y+4)=y+8 xy+4x=y+8 xy=y+8-4x xy-y=8-4x y(x-1)=8-4x y=(8-4x)/(x-1) y=(-4x+8)/(x-1) So the answer is \[\Large y = \frac{-4x+8}{x-1}\]